# Densities Of Primes And Primitive Roots

### v10

@@ 4.3 Finite Summation Kernels And Gaussian Sums
(i) Use the geometric series to compute this simple exponential sum as
$\sum_{n\leq p-1}\omega^{tn}$ $=$ $\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}\leq\bigg{|}\frac{\omega^{t}-\omega^{tp}}{1-\omega^{t}}\bigg{|}\leq\bigg{|}\frac{2}{\sin(\pi t/p)}\bigg{|}\leq\frac{2p}{\pi t}.$ (4.28)
This follows from $t\in[1,p-1]$, the sine function $\sin(\pi t/p)\neq 0$ is well defined and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$.
(ii) Using last expression (i) this becomes
$\sum_{1\leq t\leq p-1}\bigg{|}\sum_{n\leq p-1}\omega^{tn}\bigg{|}$ $\leq$ $\sum_{1\leq t\leq p-1}\frac{2p}{\pi t}\leq 2p\log p.$ (4.29)
Lemma 4.3.
  1. (i)
    $\Bigg{|}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\Bigg{|}\ll\frac{p^{1+\delta}}{\pi r(d,t)},$
    where $r(d,t)\equiv dt\bmod p$ is a permutation of $t\in\mathbb{F}_{p}$.
  2. (ii)
    $\sum_{1\leq t\leq p-1}\Bigg{|}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\omega^{tn}\Bigg{|}\ll p^{1+\delta}\log p,$
    where $\delta>0$ is a small number.
Proof.
(i) Use the inclusion exclusion principle to rewrite the exponential sum as
$=$ $\sum_{d\mid p-1}\mu(d)\sum_{m\leq(p-1)/d}\omega^{dtm}$
Now, the parameters are $p$ prime, $\omega=e^{i2\pi/p}$, the integers $t\in[1,p-1]$ and $d\leq p-1<p$. In addition, the map $t\longrightarrow dt$ is a permutation in the finite field $\mathbb{F}_{p}$ for any fixed $d\neq 0$. This data implies that $\pi dt/p\neq k\pi$ with $k\in\mathbb{Z}$, so the sine function $\sin(\pi dt/p)\neq 0$ is well defined. Using standard manipulations, and $z/2\leq\sin(z)<z$ for $0<|z|<\pi/2$, the last expression becomes
$\left|\frac{\omega^{dt}-\omega^{dtp}}{1-\omega^{dt}}\right|\leq\left|\frac{2}{\sin(\pi dt/p)}\right|\leq\frac{2p}{\pi r(d,t)},$ (4.31)
where $r(d,t)\equiv dt\bmod p$. Finally, the upper bound is
$\left|\sum_{d\mid p-1}\mu(d)\frac{\omega^{dt}-\omega^{dt((p-1)/d+1)}}{1-\omega^{dt}}\right|$ $\leq$ $\sum_{d\mid p-1}\left|\frac{2}{\sin(\pi dt/p)}\right|$
$\leq$ $\sum_{d\mid p-1}\frac{2p}{\pi r(d,t)}$
$\ll$ $\frac{p^{1+\delta}}{\pi r(d,t)},$
where $\sum_{d\mid p-1}1=d(p-1)\ll p^{\delta}$ is the number of divisor in p1p-1 and $\delta>0$ is a small number.
(ii) For each $d\mid p-1$, the map $t\longrightarrow z\equiv dt\bmod p$ is a permutation in the finite field $\mathbb{F}_{p}$. Thus,
$\leq$ $\sum_{d\mid p-1,}\sum_{1\leq z\leq p-1}\frac{2p}{\pi z}$
$\ll$ $p^{1+\delta}\log p,$
Lemma 4.4.
(Gauss sums) Let $p\geq 2$ and $q$ be large primes. Let $\chi(t)=e^{i2\pi t/q}$ and $\psi(t)=e^{i2\pi\tau^{t}/p}$ be a pair of characters. Then, the Gaussian sum has the upper bound
$\left|\sum_{1\leq t\leq q-1}\chi(t)\psi(t)\right|\leq 2q^{1/2}\log q.$ (4.34)
@@ 6 Estimate For The Error Term
The upper bounds for exponential sums over subsets of elements in finite rings $\mathbb{Z}/m\mathbb{Z}$ or finite fields $\mathbb{F}_{p}$ studied in Lemma 4.1 and Theorem 4.2 are used to estimate the error term $E(x)$ in the proof of Theorem 1.1.
@@
$\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\sum_{0<k\leq p-1}\psi\left((\tau^{n}-u)k\right)\ll\frac{x^{1-\varepsilon}}{\log x}$ (6.1)
for all sufficiently large numbers $x\geq 1$ and an arbitrarily small number $\varepsilon<1/16$.
@@
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\left(\sum_{0<k\leq p-1}e^{-i2\pi\frac{uk}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{k\tau^{n}}{p}}\right)$
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\left(-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{\tau^{n}}{p}}+O(p^{1/2+\delta}\log^{2}p)\right).$
The last line follows from Lemma 4.1. By Theorem 4.2, the inner sum has the upper bound
$\bigg{|}\widehat{V(1,u)}\bigg{|}$ $=$ $\left|-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{\tau^{n}}{p}}+O(p^{1/2+\delta}\log^{2}p)\right|$
$\ll$ $p^{1-\varepsilon},$
where $\varepsilon<1/2$ is an arbitrarily small number. Replacing the estimate (6) back into (6) and completing the finite sum return
$|E(x)|$ $\ll$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\cdot p^{1-\varepsilon}$
$\ll$ $\frac{1}{x^{\varepsilon}}\sum_{x\leq p\leq 2x}1$
$\ll$ $\frac{x^{1-\varepsilon}}{\log x},$
where the number of primes in the short interval $[x,2x]$ is $\pi(2x)-\pi(x)\leq 2x/\log x$. ∎
@@ 7 Densities Of Primes And Primitive Roots
$=$ $\delta(u)\left(\text{li}(2x)-\text{li}(x)\right)+O\left(\frac{x}{\log^{B}x}\right)+O\left(\frac{x^{1-\varepsilon}}{\log x}\right)$
$=$ $\delta(u)\frac{x}{\log x}+O\left(\frac{x}{\log^{B}x}\right)$
@@ 8 Application To Repeated Decimals
with $x_{i}\in\{0,1,2,\ldots,9\}$, was investigated by Gauss and earlier authors centuries ago, see [4] for a historical account, and [32] for recent developments. As discussed in Articles 14-18 in [16], the period, denoted by $\operatorname{ord}_{p}(10)=d\geq 1$, is a divisor of p1p-1. The problem of computing the densities for the subsets of primes for which the repeating decimals have very large periods such as d=(p1)/2d=(p-1)/2, and d=p1d=p-1, is a recent problem. This note considers the following result.
@@
There are infinitely many primes $p\geq 7$ with maximal repeating decimal $1/p=0.\overline{x_{p-2}x_{p-3}\cdots x_{1}x_{0}}$, where $0\leq x_{i}\leq 9$. Moreover, the counting function for these primes satisfies the lower bound
$\pi_{10}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(10)=p-1\right\}=\delta(10)\frac{x}{\log x}+O\left(\frac{x}{\log^{2}x}\right),$ (8.2)
@@
The proof of Theorem 8.1 is a corollary of the more general result in Theorem 1.1 presented in Section 7. This analysis generalizes to repeating $\ell$-adic expansions $1/p=0.\overline{x_{d-1}x_{d-2}\cdots x_{1}x_{0}}$, where $0\leq x_{i}\leq\ell-1$, in any numbers system with nonsquare integer base $\ell\geq 2$.
@@
are also important in the calculation of the period modulo $p^{2}$, confer [38, p. 333], [10] for other details.
@@
$\pi_{10}(x)=\sum_{p\leq x}\Psi(10)=\delta(10)\operatorname{li}(x)+O\left(\frac{x}{\log^{B}x}\right),$ (8.5)

### v12

@@ 5.2 FFT of Power Exponential Sum with Relatively Prime Index
The second line in the estimation of the upper bound in (5.2) follows from Theorem 5.2 and the fourth line follows from Theorem 5.1:
[…]
$=$ $\left|-\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{\tau^{n}}{p}}+O(p^{1/2+\delta}\log^{2}p)\right|$
$\ll$ $p^{1/2+\delta}(\log p)^{2},$
where $\delta>0$ is a small number. ∎
@@ 7 Estimate For The Error Term
The upper bounds for exponential sums over subsets of elements in finite rings $\mathbb{Z}/m\mathbb{Z}$ or finite fields $\mathbb{F}_{p}$ studied in Section 5 are used to estimate the error term $E(x)$ in the proof of Theorem 1.1.
@@
$\sum_{x\leq p\leq 2x}\frac{1}{p}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}\sum_{0<k\leq p-1}\psi\left((\tau^{n}-u)k\right)\ll x^{1/2+\delta}\log x$ (7.1)
for all sufficiently large numbers $x\geq 1$ and an arbitrarily small number $\delta>0$.
@@
$=$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\left(\sum_{0<k\leq p-1}e^{-i2\pi\frac{uk}{p}}\sum_{\begin{subarray}{c}1\leq n\leq p-1\\ \gcd(n,p-1)=1\end{subarray}}e^{i2\pi\frac{k\tau^{n}}{p}}\right).$
Applying Lemma 5.1 to the inner sum and completing the upper bound of the finite sum return
$|E(x)|$ $\ll$ $\sum_{x\leq p\leq 2x}\frac{1}{p}\cdot p^{1/2+\delta}(\log p)^{2}$
$\ll$ $\frac{(\log x)^{2}}{x^{1/2-\delta}}\sum_{x\leq p\leq 2x}1$
$\ll$ $x^{1/2+\delta}\log x,$
where the number of primes in the short interval $[x,2x]$ is $\pi(2x)-\pi(x)\leq 2x/\log x$.
@@ 8 Densities Of Primes And Primitive Roots
$=$ $\delta(u)\left(\text{li}(2x)-\text{li}(x)\right)+O\left(\frac{x}{(\log x)^{B}}\right)+O\left(x^{1/2+\delta}\log x\right)$
$=$ $\delta(u)\frac{x}{\log x}+O\left(\frac{x}{(\log x)^{B}}\right)$
@@ 9 Application To Repeated Decimals
with $x_{i}\in\{0,1,2,\ldots,9\}$, was investigated by Gauss and earlier authors centuries ago, see [4] for a historical account, and [32] for recent developments. As discussed in [16, Articles 14-18], the period, denoted by $\operatorname{ord}_{p}(10)=d\geq 1$, is a divisor of p1p-1. The problem of computing the densities for the subsets of primes for which the repeating decimals have very large periods such as d=(p1)/2d=(p-1)/2, and d=p1d=p-1, is a recent problem. This note considers the following result.
@@
There are infinitely many primes $p\geq 7$ with maximal repeating decimal
$1/p=0.\overline{x_{p-2}x_{p-3}\cdots x_{1}x_{0}},$ (9.2)
where $0\leq x_{i}\leq 9$. Moreover, the counting function for these primes satisfies the lower bound
$\pi_{10}(x)=\#\left\{p\leq x:\operatorname{ord}_{p}(10)=p-1\right\}=\delta(10)\frac{x}{\log x}+O\left(\frac{x}{(\log x)^{2}}\right),$ (9.3)
@@
The proof of Theorem 9.1 is a corollary of the more general result in Theorem 1.1 presented in Section 8. This analysis generalizes to repeating $\ell$-adic expansions
$1/p=0.\overline{x_{d-1}x_{d-2}\cdots x_{1}x_{0}},$ (9.4)
where $0\leq x_{i}\leq\ell-1$, in any numbers system with nonsquare integer base $\ell\geq 2$.
@@
are also important in the calculation of the period modulo $p^{2}$, confer [RP96, p. 333], [10] for other details.
@@
$\pi_{10}(x)=\sum_{p\leq x}\Psi(10)=\delta(10)\operatorname{li}(x)+O\left(\frac{x}{(\log x)^{B}}\right),$ (9.7)